f(x)=4x+8
g(x)=2x-12
(f+g)(x)
(f-g)(x)
(f*g)(x)

Answer by TimothyLamb(4379)

(f+g)(x) = 4x + 8 + 2x - 12
(f+g)(x) = 6x - 4
---
(f-g)(x) = 4x + 8 - 2x + 12
(f-g)(x) = 2x + 20
---
(f*g)(x) = (4x + 8)(2x - 12)
(f*g)(x) = 8xx - 48x + 16x - 96
(f*g)(x) = 8xx - 32x - 96
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---